Explain the terms stress and strain.

Grade:11

1 Answers

SAGAR SINGH - IIT DELHI
12 years ago

Stress

Stress is the ratio of applied force F and cross section A, defined as "force per area".

Direct Stress or Normal Stress

Stress normal to the plane is usually denoted "normal stress" and can be expressed as

σ = Fn / A         (1)

where

σ = normal stress ((Pa) N/m2, psi)

Fn = normal component force (N, lbf)

A = area (m2, in2)

Shear Stress

Stress parallel to the plane is usually denoted "shear stress" and can be expressed as

τ = Fp / A         (2)

where

τ = shear stress ((Pa) N/m2, psi)

Fp = parallel component force (N, lbf)

A = area (m2, in2)

Strain

Strain is defined as "deformation of a solid due to stress" and can be expressed as

ε = dl / lo = σ / E         (3)

where

dl = change of length (m, in)

lo = initial length (m, in)

ε = unitless measure of engineering strain

E = Young's modulus (Modulus of Elasticity) (Pa, psi)